# Upper bounds on the genus of Albanese fibrations

### v1

@@ 3. Upper bound on hyperelliptic Albanese fibrations with $p_{g}=q=1$
The aim in this section is to prove Theorem 1.2. So we always assume in this section that $S$ is a minimal irregular surface of general type with $p_{g}=q=1$ and its Albanese fibration $f:S\rightarrow B$ is hyperelliptic. In this case, $g(B)=q=1$ and $f$ must be locally non-trivial. By (2.4), the relative invariants equal the corresponding invariants of $S$. We use singularity indices $s_{i}$’s introduced in Section 2.2. Since it is unknown whether $s_{2}$ is nonnegative, we divide the proof into two cases, depending on whether $s_{2}<0$ or $s_{2}\geq 0$. The proof of Theorem 1.2 will be completed in Section 3, Section 3 and Section 3.
@@
Let $f:\,S\to B$ be as in Theorem 1.2. If $s_{2}<0$, then $g\leq 5$.
@@
$\frac{3}{2}\leq\frac{(n+1)^{2}}{3n}\leq\chi_{\rm top}(S)-\frac{1}{3}K_{S}^{2}=\frac{4}{3}(9\chi-K_{S}^{2}).$
@@
Come back to the proof of Section 3. By Theorem 2.1, we have
@@
=\,= $-\frac{g-2}{2(g-1)}s_{2}+\frac{g^{2}+2g-5}{2(g-1)}s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}\left(\frac{2k(g-k)}{g-1}-1\right)s_{2k+1}+\sum_{k=3}^{[\frac{g+1}{2}]}\left(\frac{k(g-k+1)}{g-1}-2\right)s_{2k}$
[…]
@@
In the rest part of this section, we assume that $s_{2}\geq 0$. We first claim that
@@
Let $f:\,S\to B$ be as in Theorem 1.2. Suppose that $s_{2}\geq 0$. Then
@@
Hence $g<6$. This shows that $s_{g+2}=0$ if $g\geq 6$. The rest two inequalities can be proved similarly using (2.6), and left to the readers. ∎
@@
Let $f:\,S\to B$ be as in Theorem 1.2. If $g\geq 13$, then there is only one possible case for $(K_{S}^{2},g)$: $K_{S}^{2}=8,g=14$. In particular, $g\leq 14$.
@@
Since $g\geq 13$ and each $s_{i}$ is a nonnegative integer, it follows that $s_{3}=s_{4}=s_{6}=s_{8}=0$, and hence $11-\frac{g}{2}=\frac{g}{4}s_{2}$, i.e., $s_{2}+2=\frac{44}{g}$, which implies that $s_{2}=0$ and $g=22$. Substitute into the formulas of Theorem 2.1, one obtains that $K_{S}^{2}=K_{f}^{2}=10>9=9\chi(\mathcal{O}_{S})$, which is a contradiction. (2) We show that if $s_{8}\neq 0$, then
@@
In fact, By (1), we already prove $s_{10}=0$. According to (3.2), $s_{8}=1$ and
@@
Since $g\geq 13$, it follows that $s_{3}=s_{4}=s_{6}=0$, and $gs_{2}=28$. Hence either $g=28,s_{2}=1$ or $g=14,s_{2}=2$. The first case is impossible; otherwise, by (3.3) one has $n=\frac{1}{2}$, which is a contradiction since $n$ is an integer. Finally, $K_{S}^{2}=8$ follows from the formula in Theorem 2.1. (3) We show that $s_{6}=0$. Otherwise we would have $s_{6}=1$ again by (3.2). Moreover, by the arguments (1) and (2) above, we have $s_{10}=s_{8}=0$. Moreover, we have
@@
Hence $s_{3}=s_{4}=0$, $s_{2}=3$ and $g=16$. Substitute into (3.3), one obtains that $n=\frac{1}{2}$, which gives a contradiction as $n$ is an integer. (4) We show that $s_{3}=s_{4}=0$. Otherwise by (2), $s_{8}=0$. Combining with (1) and (2) above, we have $s_{6}=s_{8}=s_{10}=0$, and hence by (3.2)
@@
Hence $s_{3}+s_{4}\leq 2$. If $s_{3}+s_{4}=2$, then $gs_{2}=12$, which is impossible since $g\geq 13$. If $s_{3}+s_{4}=1$, then by (3.3) one gets $gn=4$, which is impossible since $g\geq 13$ and $n\geq 1$. In conclusion, if $g\geq 13$, then except the possible case $(K_{S}^{2},g)=(8,14)$, one has $s_{i}=0$ for all $i>2$. Hence by (3.3) it holds $gn=2$, which is impossible since $n$ is an integer. This completes the proof. ∎
@@
Let $f:\,S\to B$ be as in Theorem 1.2.
@@
Note that all the $s_{i}$’s are non-negative integers. It is not difficult to show that $s_{6}=s_{8}=s_{10}=s_{12}=0$. Thus $253s_{2}+11(s_{3}+s_{4})=$ 25 . It follows that $s_{2}=1$ and $s_{3}+s_{4}=2$. By Theorem 2.1 one has $36n=\frac{gn}{2}=1+s_{3}+s_{4}=$ 3, which is impossible since $n$ is an integer. (2) Suppose that $g=11$. Similar as above, by (3.1) together with Theorem 2.1,
@@
One checks again this is impossible since all $s_{i}$’s are non-negative integers. This completes the proof. ∎
@@
Let $f:\,S\to B$ be as in Theorem 1.2. By a similar argument as above, one can show that $g\neq 9$. More precisely, the list of all possibilities for $(K_{S}^{2},g)$ is as follow:

### v3

@@ 2.2. Invariants of hyperelliptic fibrations
3. Quadratic upper bounds on hyperelliptic Albanese fibrations
The main purpose of this section is to prove Theorem 1.1. We will prove the following upper bounds (using the relative invariants) on the genus $g$ for any not locally trivial hyperelliptic fibration, not necessarily the Albanese fibration. We remark once more that the relative invariants of $f$ equal the corresponding invariants of the surface $S$ (cf. (2.5)) if the base $B$ is an elliptic curve.
Theorem 3.1.
Let $f:\,S\to B$ be a relatively minimal hyperelliptic fibration of genus $g\geq 2$. Suppose that $f$ is not locally trivial.
  1. (1)
    If $\lambda_{f}\leq 4$, then
    $g\leq\frac{4\chi_{f}+4}{2+(4\chi_{f}-K_{f}^{2})}\leq 2\chi_{f}+2.$
  2. (2)
    If $\lambda_{f}>4$, then
    $g\leq\left\{\begin{aligned} &2\chi_{f}+1,&&\text{if $g(B)=0$};\\ &\frac{25}{4}\chi_{f}^{2}+\frac{19}{2}\chi_{f}+2,&\quad&\text{if $g(B)=1$};\\ &16\chi_{f}^{2}+14\chi_{f}+2,&&\text{if $g(B)\geq 2$}.\end{aligned}\right.$
We first prove Theorem 1.1 based on Theorem 3.1.
Proof of Theorem 1.1.
Let $f:\,S\to B$ be the Albanese fibration of $S$ and let $g$ be the genus of a general fiber of $f$. Since $g(B)=q(S)=1$, $f$ must be not locally trivial; otherwise $\chi(\mathcal{O}_{S})=\chi_{f}=0$ which is absurd. Thus by (2.5) together with Theorem 3.1 (2), one obtains (1.1). ∎
[…]
(1) If $\lambda_{f}=4$ and $g(B)=1$, then the upper bound $g\leq 2\chi_{f}+2$ has already been obtained by Ishida [Ish05].
(2) We will construct examples in Section 5 showing that the quadratic upper bounds can not be improved into a linear one.
The proof of Theorem 3.1 relies on the following technical result.
Theorem 3.2.
Let $f:\,S\to B$ be a relatively minimal hyperelliptic fibration of genus $g\geq 2$, and $n=\frac{R^{2}}{4(g+1)}$ be an integer as in (2.2). Suppose that $f$ is not locally trivial with slope $\lambda_{f}>4$ and $g(B)\geq 1$. Then
$1\leq n\leq\Big{(}\frac{\lambda_{f}-4}{2}+\frac{\lambda_{f}}{g-1}\Big{)}\chi_{f},$ (3.1)
and
$g\leq\frac{(\lambda_{f}-4)^{2}}{4n}\chi_{f}^{2}+(\lambda_{f}-4+\frac{\lambda_{f}}{2n})\chi_{f}+n+1.$ (3.2)
Moreover, equality in (3.2) holds if and only if there exists exactly one $s_{2k}=1$ for some $k\geq 2$ and $s_{i}=0$ for all $i\geq 3,i\neq 2k$.
We first prove Theorem 3.1 based on Theorem 3.2 in the following, and then prove Theorem 3.2.
Proof of Theorem 3.1.
(1). Note that if $s_{i}=0$ for all $i\geq 3$, then by Theorem 2.1 we get $\chi_{f}=\frac{1}{2}gn$ and consequently $g=\frac{2\chi_{f}}{n}\leq 2\chi_{f}<2\chi_{f}+2$. So we can assume that $s_{i}\neq 0$ for some $i\geq 3$.
Since $\lambda_{f}\leq 4$, i.e., $K_{f}^{2}\leq 4\chi_{f}$, it follows that $t:=4\chi_{f}-K_{f}^{2}\geq 0$. By Theorem 2.1 one gets
$\chi_{f}-\frac{gt}{4}$ $\,=\frac{g}{4}\left(K_{f}^{2}-\frac{4(g-1)}{g}\chi_{f}\right)$
$\,=\frac{g-2}{2}.$
Thus $g\leq\frac{4\chi_{f}+4}{2+t}\leq 2\chi_{f}+2$ as required.
(2). If $g(B)=0$, then
$\chi_{f}=\chi(\mathcal{O}_{S})+(g-1)=p_{g}-q+g\geq\frac{g-1}{2}.$
The inequality above follows from the inequality $q\leq\frac{g+1}{2}$ by Xiao [Xia87b] and the non-negativity of $p_{g}$. Hence $g\leq 2\chi_{f}+1$ as required.
We consider next the case when $g(B)\geq 1$. By Theorem 3.2 above, we have
$g\leq\frac{(\lambda_{f}-4)^{2}}{4n}\chi_{f}^{2}+(\lambda_{f}-4+\frac{\lambda_{f}}{2n})\chi_{f}+n+1,\qquad n\leq\big{(}\frac{\lambda_{f}-4}{2}+\frac{\lambda_{f}}{g-1}\big{)}\chi_{f}.$ (3.3)
If $g(B)=1$, then $\lambda_{f}\leq 9$ by (2.6), and hence
$g\leq\max_{1\leq n\leq\big{(}\frac{\lambda_{f}-4}{2}+\frac{\lambda_{f}}{g-1}\big{)}\chi_{f}}\left\{g(n):=\frac{25}{4n}\chi_{f}^{2}+\Big{(}5+\frac{9}{2n}\Big{)}\chi_{f}+n+1\right\}.$
Suppose that
$g>\frac{25}{4}\chi_{f}^{2}+\frac{19}{2}\chi_{f}+2=g(1).$ (3.4)
Since $\chi_{f}\geq 1$, we have $g\geq 18$ and thus $n\leq 3\chi_{f}$. Note that as $n$ increases, $g(n)$ first decreases and then increases. Hence we have
$g\leq\max_{1\leq n\leq\big{(}\frac{\lambda_{f}-4}{2}+\frac{\lambda_{f}}{g-1}\big{)}\chi_{f}}\{g(n)\}\leq\max\{g(1),g(3\chi_{f})\}=\max\{g(1),\frac{121}{12}\chi_{f}+\frac{5}{2}\}=g(1),$
which contradicts to (3.4).
If $g(B)=2$, the argument is almost the same as the case when $g(B)=1$, except that we have to replace (2.6) by (2.4); the details are omitted here. ∎
We now prove the technical result Theorem 3.2.
Proof of Theorem 3.2.
We first prove (3.2). By Theorem 2.1, we have
$(\lambda_{f}-4)\chi_{f}+2n=K_{f}^{2}-4\chi_{f}+2n=(2g+2)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}2(k-1)s_{2k}.$ (3.5)
Hence
$(\lambda_{f}-4)^{2}\chi_{f}^{2}+(4n+3)(\lambda_{f}-4)\chi_{f}+4n^{2}+6n$ (3.6)
$=$ $\big{(}(\lambda_{f}-4)\chi_{f}+2n\big{)}^{2}+3\big{(}(\lambda_{f}-4)\chi_{f}+2n\big{)}$
$=$ $\Big{(}(2g+2)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}2(k-1)s_{2k}\Big{)}^{2}$
$+3\Big{(}(2g+2)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}2(k-1)s_{2k}\Big{)}$
$\geq$ $(g+1)(4g+10)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(16k^{2}+4k-2)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}(4k^{2}-2k-2)s_{2k}$
$\geq$ $(g+1)(2g+4)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}(4k^{2}-2k-2)s_{2k}.$
On the other hand, by Theorem 2.1, we have
$\chi_{f}$ $\,=\frac{gs_{2}+(g^{2}-2g-1)s_{g+2}}{4(2g+1)}+\sum_{k=1}^{[\frac{g}{2}]}\frac{k(g-k)}{2g+1}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{k(g-k+1)}{2(2g+1)}s_{2k}$
$\,=\frac{g}{4(2g+1)}\Big{(}e_{f}-\sum_{k=1}^{[\frac{g}{2}]}s_{2k+1}-\sum_{k=2}^{[\frac{g+1}{2}]}2s_{2k}\Big{)}+\frac{(g^{2}-1)s_{g+2}}{4(2g+1)}+\sum_{k=1}^{[\frac{g}{2}]}\frac{k(g-k)}{2g+1}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{k(g-k+1)}{2(2g+1)}s_{2k}$
$\,=\frac{1}{8}\Big{(}e_{f}+(2g+2)s_{g+2}\Big{)}+\sum_{k=1}^{[\frac{g}{2}]}\big{(}\frac{k}{2}-\frac{1}{8}\big{)}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\big{(}\frac{k}{4}-\frac{1}{4}\big{)}s_{2k}$
$\quad-\frac{1}{2g+1}\left(\frac{1}{8}\Big{(}e_{f}+(g+1)(2g+4)s_{g+2}\Big{)}+\sum_{k=1}^{[\frac{g}{2}]}\big{(}\frac{2k^{2}+k}{2}-\frac{1}{8}\big{)}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\big{(}\frac{2k^{2}-k}{4}-\frac{1}{4}\big{)}s_{2k}\right)$
Combining this with (3.5), we get
$2n$ $\,=e_{f}+(2g+2)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}2(k-1)s_{2k}-8\chi_{f}$
$\,=\frac{1}{2g+1}\bigg{(}e_{f}+(g+1)(2g+4)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}(4k^{2}-2k-2)s_{2k}\bigg{)}.$ (3.7)
Note that the left hand side of the above equality is an integer and that the right hand side of the above equality is positive. Hence we have
$e_{f}+(g+1)(2g+4)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}(4k^{2}-2k-2)s_{2k}=2n(2g+1).$ (3.8)
Now combining (3.6) and (3.8), we get
$\,(\lambda_{f}-4)^{2}\chi_{f}^{2}+\big{(}4n(\lambda_{f}-4)+2\lambda_{f}\big{)}\chi_{f}+4n^{2}+6n$
$\,=e_{f}+(\lambda_{f}-4)^{2}\chi_{f}^{2}+(4n+3)(\lambda_{f}-4)\chi_{f}+4n^{2}+6n$
$\,\geq e_{f}+(g+1)(2g+4)s_{g+2}+\sum_{k=1}^{[\frac{g+1}{2}]}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}(4k^{2}-2k-2)s_{2k}$
$\,=2n(2g+1).$ (3.9)
Hence we get
$g\leq\frac{(\lambda_{f}-4)^{2}}{4n}\chi_{f}^{2}+\big{(}\lambda_{f}-4+\frac{\lambda_{f}}{2n}\big{)}\chi_{f}+n+1.$
Finally, the equality in (3.2) holds if and only if equality in (3.6) holds, and if and only if there exists exactly one $s_{2k}=1$ for some $k\geq 2$ and $s_{i}=0$ for all $i\geq 3,i\neq 2k$.
Now we prove (3.1). According to (2.3) and (3.7), we get
$\frac{(\lambda_{f}-4)g+4}{g}\chi_{f}$ $\geq\frac{g-1}{2g}\left((2g+2)s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}(4k-1)s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}2(k-1)s_{2k}\right)$
$\geq\frac{g-1}{2g}\big{(}(\lambda_{f}-4)\chi_{f}+2n\big{)}.$
This proves (3.1). ∎
[…]
We will construct examples reaching the equality in (3.2); see Section 5.
If $\chi_{f}$ is large, one can obtain a better upper bound on $g$ as follows.
[…]
Let $f:\,S\to B$ be a relatively minimal hyperelliptic fibration of genus $g\geq 2$, and $n=\frac{R^{2}}{4(g+1)}$ be an integer as in (2.2). Suppose that $f$ is not locally trivial with slope $\lambda_{f}>4$ and $g(B)\geq 1$. Set $g(n):=\frac{(\lambda_{f}-4)^{2}}{4n}\chi_{f}^{2}+\big{(}\lambda_{f}-4+\frac{\lambda_{f}}{2n}\big{)}\chi_{f}+n+1$. If $g\geq 25$ and $\chi_{f}\geq 4$, then we have
$g(1)\geq g(2)\geq\max_{1\leq n\leq\big{(}\frac{\lambda_{f}-4}{2}+\frac{\lambda_{f}}{g-1}\big{)}\chi_{f}}\{g(n)\}.$
Proof.
(1) Since $\chi_{f}\geq 1$ and $\lambda_{f}\geq 4$, we have $\frac{\lambda_{f}}{2}\chi_{f}+1\geq\frac{\lambda_{f}}{4}\chi_{f}+2$ and thus $g(1)\geq g(2)$. So we only need to prove
$g(2)\geq\max\limits_{1\leq n\leq\big{(}\frac{\lambda_{f}-4}{2}+\frac{\lambda_{f}}{g-1}\big{)}\chi_{f}}\{g(n)\}.$
Since $n\leq\big{(}\frac{\lambda_{f}-4}{2}+\frac{\lambda_{f}}{g-1}\big{)}\chi_{f}$, $\lambda_{f}<12$ (see (2.4)) and we have assumed $g\geq 25$, we get $n<\frac{\lambda_{f}-3}{2}\chi_{f}$. Note that as $n$ increases, $g(n)$ first decreases and then increases. Hence we only need to show $g(2)\geq g\big{(}\frac{\lambda_{f}-3}{2}\chi_{f}\big{)}$.
$g(2)-g\big{(}\frac{\lambda_{f}-3}{2}\chi_{f}\big{)}=\Big{(}\frac{\chi_{f}}{8}-\frac{1}{2(\lambda_{f}-3)}\Big{)}(\lambda_{f}-4)^{2}\chi_{f}+\frac{6-\lambda_{f}}{4}\chi_{f}+2-\frac{\lambda_{f}}{\lambda_{f}-3}.$
Since we have assumed $\lambda_{f}\geq 4$ and $\chi_{f}\geq 4$, we have $\frac{\chi_{f}}{8}\geq\frac{1}{2}\geq\frac{1}{2(\lambda_{f}-3)}$.
(i) If $\lambda_{f}\leq 6$, then we have
$g(2)-g\big{(}\frac{\lambda_{f}-3}{2}\chi_{f}\big{)}\geq 6-\lambda_{f}+2-\frac{\lambda_{f}}{\lambda_{f}-3}=4-\left((\lambda_{f}-3)+\frac{3}{\lambda_{f}-3}\right)\geq 0.$
(ii) If $\lambda_{f}>6$, then we have $\lambda_{f}-3\geq 3$ and $2-\frac{\lambda_{f}}{\lambda_{f}-3}\geq 0$. Hence we have
$g(2)-g\big{(}\frac{\lambda_{f}-3}{2}\chi_{f}\big{)}\geq\left(\big{(}\frac{1}{2}-\frac{1}{6}\big{)}(\lambda_{f}-4)^{2}+\frac{1}{2}-\frac{\lambda_{f}-4}{4}\right)\chi_{f}>0.$
Therefore if $g\geq 25$ and $\chi_{f}\geq 4$, we always have $g(2)\geq g\big{(}\frac{\lambda_{f}-3}{2}\chi_{f}\big{)}$. ∎
[…]
Let $f:\,S\to B$ be a relatively minimal hyperelliptic fibration of genus $g\geq 2$. Suppose that $f$ is not locally trivial with slope $\lambda_{f}\geq 4$ and $g(B)\geq 1$. If $g\geq 25$ and $\chi_{f}\geq 4$, then we have
$g\leq\left\{\begin{aligned} &\frac{(\lambda_{f}-4)^{2}}{4}\chi_{f}^{2}+(\frac{3}{2}\lambda_{f}-4)\chi_{f}+2,&\quad&\text{if $g$ is even};\\ &\frac{(\lambda_{f}-4)^{2}}{8}\chi_{f}^{2}+(\frac{5}{4}\lambda_{f}-4)\chi_{f}+3,&&\text{if $g$ is odd}.\end{aligned}\right.$ (3.10)
Moreover, equality in (3.10) holds if and only if there exists exactly one $s_{2k}=1$ for some $k\geq 2$ and $s_{i}=0$ for all $i\geq 3,i\neq 2k$.
Proof.
By Theorem 3.2 and Section 3, we only need to prove $n\geq 2$ if $g$ is odd. Here we use notations as in Section 2.2. Recall that $R\sim-(g+1)K_{P/B}+n\Gamma$ is an even divisor. If $g$ is odd, then $-(g+1)K_{P/B}$ is an even divisor and thus $n\Gamma$ is also an even divisor. Since the Néron-Severi group $\text{Num}(P)$ is generated by a section and a fibre $\Gamma$ of $h:\,P\rightarrow B$, we see $n$ is even. Since $n$ is a positive integer by (3.7), we get $n\geq 2$. ∎
Since we have $\lambda_{f}\leq 9$ if $g(B)=1$ (see (2.6)) and $\lambda_{f}<12$ if $g(B)\geq 2$ (see (2.4)), we get the following.
[…]
Under the same assumptions as in Section 3 the following hold.
(1) If $g(B)=1$, we have
$g\leq\left\{\begin{aligned} &\frac{25}{4}\chi_{f}^{2}+\frac{19}{2}\chi_{f}+2,&\quad&\text{if $g$ is even};\\ &\frac{25}{8}\chi_{f}^{2}+\frac{29}{4}\chi_{f}+3,&&\text{if $g$ is odd}.\end{aligned}\right.$ (3.11)
(1) If $g(B)\geq 2$, we have
$g\leq\left\{\begin{aligned} &16\chi_{f}^{2}+14\chi_{f}+2,&\quad&\text{if $g$ is even};\\ &8\chi_{f}^{2}+11\chi_{f}+3,&&\text{if $g$ is odd}.\end{aligned}\right.$ (3.12)
@@ 4. Upper bound on hyperelliptic Albanese fibrations with $p_{g}=q=1$
The aim in this section is to prove Theorem 1.2. So we always assume in this section that $S$ is a minimal irregular surface of general type with $p_{g}=q=1$ and its Albanese fibration $f:S\rightarrow B$ is hyperelliptic. In this case, $g(B)=q=1$ and $f$ is not locally trivial otherwise $0=\chi_{f}=\chi(\mathcal{O}_{S})$. By (2.5), the relative invariants equal the corresponding invariants of $S$. We use the singularity indices $s_{i}$’s introduced in Section 2.2. Since it is unknown whether $s_{2}$ is nonnegative, we divide the proof into two cases, depending on whether $s_{2}<0$ or $s_{2}\geq 0$. The proof of Theorem 1.2 will be completed in Section 4, Section 4 and Section 4.
@@
Let $S$ be a minimal surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ of $S$ is hyperelliptic of genus $g$. If $s_{2}<0$, then $g\leq 5$.
@@
$\frac{3}{2}\leq\frac{(n+1)^{2}}{3n}\leq e(S)-\frac{1}{3}K_{S}^{2}=\frac{4}{3}(9\chi-K_{S}^{2}).$
@@
Come back to the proof of Section 4. Since $s_{2}$ is an integer and we have assumed that $s_{2}<0$, we see that $s_{2}\leq-1$. By Theorem 2.1, we have
@@
$=$ $-\frac{g-2}{2(g-1)}s_{2}+\frac{g^{2}+2g-5}{2(g-1)}s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}\left(\frac{2k(g-k)}{g-1}-1\right)s_{2k+1}+\sum_{k=3}^{[\frac{g+1}{2}]}\left(\frac{k(g-k+1)}{g-1}-2\right)s_{2k}$
$\geq$ $\frac{g-2}{2(g-1)}.$
@@
In the remaining part of this section, we assume that $s_{2}\geq 0$. We first claim that
@@
Let $S$ be a minimal surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ of $S$ is hyperelliptic of genus $g$. Suppose that $s_{2}\geq 0$. Then
@@
Hence $g<6$. This shows that $s_{g+2}=0$ if $g\geq 6$. The other two inequalities can be proved similarly using (2.7), and are left to the readers. ∎
@@
Let $S$ be a minimal surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ of $S$ is hyperelliptic of genus $g$. If $g\geq 13$, then there is only one possible case for $(K_{S}^{2},g)$: $K_{S}^{2}=8,g=14$. In particular, $g\leq 14$.
@@
Since $g\geq 13$ and each $s_{i}$ is a nonnegative integer, it follows that $s_{3}=s_{4}=s_{6}=s_{8}=0$, and hence $11-\frac{g}{2}=\frac{g}{4}s_{2}$, i.e., $s_{2}+2=\frac{44}{g}$, which implies that $s_{2}=0$ and $g=22$. Substitute into the formulas of Theorem 2.1, one obtains that $K_{S}^{2}=K_{f}^{2}=10>9=9\chi(\mathcal{O}_{S})$, which is a contradiction.
(2) We show that if $s_{8}\neq 0$, then
@@
In fact, by (1), we have already proven $s_{10}=0$. According to (4.2), $s_{8}=1$ and
@@
Since $g\geq 13$, it follows that $s_{3}=s_{4}=s_{6}=0$, and $gs_{2}=28$. Hence either $g=28,s_{2}=1$ or $g=14,s_{2}=2$. The first case is impossible; otherwise, by (4.3) one has $n=\frac{1}{2}$, which is a contradiction since $n$ is an integer. Finally, $K_{S}^{2}=8$ follows from the formula in Theorem 2.1.
(3) We show that $s_{6}=0$. Otherwise we would have $s_{6}=1$ again by (4.2). Moreover, by the arguments (1) and (2) above, we have $s_{10}=s_{8}=0$. Moreover, we have
@@
Hence $s_{3}=s_{4}=0$, $s_{2}=3$ and $g=16$. Substitute into (4.3), one obtains that $n=\frac{1}{2}$, which gives a contradiction as $n$ is an integer.
(4) We show that $s_{3}=s_{4}=0$. Otherwise by (2), $s_{8}=0$. Combining with (1) and (2) above, we have $s_{6}=s_{8}=s_{10}=0$, and by (4.2)
@@
Hence $s_{3}+s_{4}\leq 2$. If $s_{3}+s_{4}=2$, then $gs_{2}=12$, which is impossible since $g\geq 13$. If $s_{3}+s_{4}=1$, then by (4.3) one gets $gn=4$, which is impossible since $g\geq 13$ and $n\geq 1$.
In conclusion, if $g\geq 13$, then except the possible case $(K_{S}^{2},g)=(8,14)$, one has $s_{i}=0$ for all $i>2$. Hence by (4.3) it holds $gn=2$, which is impossible since $n$ is an integer. This completes the proof. ∎
@@
Let $S$ be a minimal surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ of $S$ is hyperelliptic of genus $g$.
@@
Note that all the $s_{i}$’s are non-negative integers. It is not difficult to show that $s_{6}=s_{8}=s_{10}=s_{12}=0$. Thus $253s_{2}+11(s_{3}+s_{4})=$ 25 . It follows that $s_{2}=1$ and $s_{3}+s_{4}=2$. By Theorem 2.1 one has $36n=\frac{gn}{2}=1+s_{3}+s_{4}=$ 3, which is impossible since $n$ is an integer.
(2) Suppose that $g=11$. Similar as above, by (4.1) together with Theorem 2.1,
@@
One checks again that this is impossible since all $s_{i}$’s are non-negative integers. This completes the proof. ∎
@@
Let $S$ be a minimal surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ of $S$ is hyperelliptic of genus $g$. By a similar argument as above, one can show that $g\neq 9$. More precisely, the list of all possibilities for $(K_{S}^{2},g)$ is as follow: